OpenStudy (anonymous):

Medal and fan 1. Maria and Ming start at the same point and drive in opposite directions. Maria drives 38 mph and Ming drives 40 mph. How far apart will they be in 2 hours? A:78 miles B:156 miles C:166 miles D:116 miles

OpenStudy (anonymous):

40*2=80 38*2=76 80+76=156 Answer: B

OpenStudy (anonymous):

can you help me with a few more

OpenStudy (anonymous):

I gotta do my algebra 2 homework first

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you for the help though